OpenStudy (anonymous):

what is the absoulte vaule of 12-2i

OpenStudy (anonymous):

\[12 + 2i\]

OpenStudy (anonymous):

2(6+i)

OpenStudy (anonymous):

Red'bone: it's not the absolute value that you are looking for, but the "magnitude". And, for "a+bi", the magnitude is |a+bi| = sqrt( a^2 + b^2 )

OpenStudy (anonymous):

caveman has it, solve it,and post ur answer,we'll check if its rite :)

OpenStudy (anonymous):

the answwer was wrong

OpenStudy (anonymous):

ok the answer is sqrt(12^2+(-2)^2)=sqrt(148)=12.17. concept of magnitude of a complex number is, wen u hav a+ib then in the argand plane(a plane like x-y,where x represents real values and y represents imaginary values) its just a point (a,b).by magnitude , u mean the length of the linesegment joining (a,b) to (0,0). so use distance formula to find it :)

OpenStudy (anonymous):

get it?